OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x

OpenStudy (anonymous):

x + 20/2 = 3x What do you think you need to do first?

OpenStudy (anonymous):

I honestly have no Idea

OpenStudy (anonymous):

Okay. First you need to multiply both sides by 2 to get rid of the 2.

OpenStudy (anonymous):

2(x+20/2) = 3x(2) Then the 2's cancel eachother out on the left side x+20=6x Then you want to get rid of the x's on the same side so you would subtract x from eachside 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

OpenStudy (texaschic101):

nicely done cydnimarie :)

OpenStudy (anonymous):

:)